Cover-to-Cover Company

Contribution Margin Income Statement

For the Year Ended December 31

1

Sales

$424,000.00

2

Variable costs:

3

Manufacturing

$233,200.00

4

Selling

21,200.00

5

Administrative

63,600.00

318,000.00

6

Contribution margin

$106,000.00

7

Fixed costs:

8

Manufacturing

$5,000.00

9

Selling

4,000.00

10

Administrative

33,400.00

42,400.00

11

Income from operations

$63,600.00

Review the contribution margin income statements for Cover-to-Cover Company and Biblio Files Company on their respective Income Statements panels. Complete the following table from the data provided in the income statements. Each company sold 84,800 units during the year.

Cover-to-Cover Company Biblio Files Company
Contribution margin ratio (percent)
Unit contribution margin
Break-even sales (units)
Break-even sales (dollars)
Answer #1

Calculate following

Cover to cover company
Contribution margin ratio 106000/424000 = 25%
Unit contribution margin 106000/84800 = 1.25 per unit
Break even sales (units) 42400/1.25 = 33920 Units
Break even sales (dollars) 42400/.25 = $169600

Note : Biblio Files Company income statement is not given